# How to Find the Exact Pareto Front for Multi-Objective MDPs?

### v1

@@ 4.3 Locality Property of the Pareto front
Let deterministic policy $\pi$ correspond to a vertex on ${\mathbb{J}}$. We denote the set of neighbors on the boundary of ${\mathbb{J}}$ to $J^{\pi}$ as $\mathrm{N}(J^{\pi},{\mathbb{J}})$, and the set of deterministic policies whose long-term returns make up $\mathrm{N}(J^{\pi},{\mathbb{J}})$ as $\mathrm{N}_{\pi}(J^{\pi},{\mathbb{J}})$. Let ${\mathbb{F}}(J^{\pi},{\mathbb{J}})$ and ${\mathbb{F}}(J^{\pi},\mathrm{Conv}(J^{\Pi_{1}}\cup\{J^{\pi}\})$ be the set of incident faces of $J^{\pi}$ on ${\mathbb{J}}$ and $\mathrm{Conv}(J^{\Pi_{1}}\cup\{J^{\pi}\}$, respectively.
@@
This theorem guarantees that we can find all neighboring vertices and faces on the boundary of ${\mathbb{J}}$ from the local convex hull $\mathrm{Conv}(J^{\Pi_{1}}\cup\{J^{\pi}\})$. By iterating through the process of constructing the convex hull of all distance-one policies and retaining the neighboring faces, we can eventually traverse the entire boundary of ${\mathbb{J}}$.
We also establish key properties of the Pareto front on a convex polytope.
@@
Given a convex polytope ${\mathbb{C}}$, let $F$ be face of ${\mathbb{C}}$ defined as the intersection of $n$ facets, i.e., $F=\cap_{i=1}^{n}F_{i}$, where $n\geq 1$ and each facet is of the form $F_{i}={\mathbb{C}}\cap\{{\bm{x}}\in\mathbb{R}^{d}:{\bm{w}}_{i}^{\top}{\bm{x}}=c_{i}\}$. Then the face $F$ lies on the Pareto front of ${\mathbb{C}}$ if there exists a linear combination of the normal vectors, weighted by $\boldsymbol{\alpha}_{i}\geq 0$ for any $i\in\{1,\cdots,n\}$, such that $\sum_{i}\boldsymbol{\alpha}_{i}[{\bm{w}}_{i}]_{j}>0$ where $[{\bm{w}}_{i}]_{j}$ is the $j$-th component of ${\bm{w}}_{i}$.
Based on Lemma 3, we can also extract the Pareto optimal faces from the local convex hull. Let ${\mathbb{F}}(J^{\pi},\mathbb{P}({\mathbb{J}}))$ be the set of incident faces of $J^{\pi}$ on the Pareto front $\mathbb{P}({\mathbb{J}})$.
@@ Appendix C Supporting Lemmas and Theorems
Let ${\bm{x}}$ be a point in a set ${\mathbb{S}}$. If there exists a vector ${\bm{w}}>0$ such that ${\bm{w}}^{\top}({\bm{x}}-{\bm{y}})\geq 0$ for all ${\bm{y}}\in{\mathbb{S}}$, then ${\bm{x}}$ lies on the Pareto front of ${\mathbb{S}}$.
@@
Assume, for the sake of contradiction, that ${\bm{x}}$ is not on the Pareto front of ${\mathbb{S}}$. This means there exists a point ${\bm{y}}\in{\mathbb{S}}$ that strictly dominates ${\bm{x}}$, i.e., ${\bm{y}}\succ{\bm{x}}$. This implies that ${\bm{y}}$ is at least as large as the corresponding objective of ${\bm{x}}$ and ${\bm{y}}$ is strictly better than ${\bm{x}}$ in at least one objective. Therefore, given ${\bm{w}}>0$, we have ${\bm{w}}^{\top}({\bm{x}}-{\bm{y}})<0$. However, this contradicts the assumption that ${\bm{w}}^{\top}({\bm{x}}-{\bm{y}})\geq 0$ for all ${\bm{y}}\in{\mathbb{S}}$. Thus, no point in ${\mathbb{S}}$ strictly dominates ${\bm{x}}$, which implies that ${\bm{x}}$ is on the Pareto front. ∎
@@
[38] The occupancy measure of any policy $\pi$ and initial distribution $\mu$ is defined as $d_{\mu}^{\pi}(s)\coloneqq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s|s_{0}\sim\mu(\cdot))$. Similarly, define $\bar{d}_{\mu}^{\pi}(s,a)\coloneqq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s,a_{t}=a|s_{0}\sim\mu(\cdot))=d_{\mu}^{\pi}(s)\pi(a|s)$. Define for any set of policies $U$, $L^{U}(\mu)\coloneqq\cup_{\pi\in U}\bar{d}_{\mu}^{\pi}$ and define $L(\mu)\coloneqq\cup_{\pi\in\Pi}\bar{d}_{\mu}^{\pi}$. Then, $L(\mu)$ is closed convex polytope and $L(\mu)=\mathrm{Conv}(L{\Pi_{D}}(\mu))$.
@@
Given initial state distribution $\mu>0$, recall the definition of ${\mathbb{J}}(\mu)=\{\pi\in\Pi|J^{\pi}(\mu)\}$ is all the achievable long-term returns for any policies belonging to $\Pi$ which can be non-stationary and stochastic policies, and the Pareto front of ${\mathbb{J}}(\mu)$ is denoted as $\mathbb{P}({\mathbb{J}}(\mu))$. Lemma 1 shows that the ${\mathbb{J}}(\mu)$ construct a convex polytope. The boundary and vertices of the convex polytope ${\mathbb{J}}(\mu)$ are denoted as ${\mathbb{B}}({\mathbb{J}}(\mu))$ and ${\mathbb{V}}({\mathbb{J}}(\mu))$, respectively. Lemma 1 also implies the vertices of ${\mathbb{J}}(\mu)$ are deterministic policies, that is, ${\mathbb{V}}({\mathbb{J}}(\mu))\subseteq\Pi_{D}$.
@@
(Restatement of Lemma 1) ${\mathbb{J}}(\mu)$ is closed convex polytope, and the vertices of ${\mathbb{J}}(\mu)$ can be achieved by deterministic policies.
@@
For any edge on ${\mathbb{B}}({\mathbb{J}}(\mu))$ with endpoints denoted as $J^{\pi_{1}}(\mu)$ and $J^{\pi_{2}}(\mu)$, any point on this edge can only be achieved by the long-term return of a convex combination of policies $\pi_{1}$ and $\pi_{2}$. Specifically, for any $\alpha\in[0,1]$, there exists a $\beta\in[0,1]$ such that
@@
We construct a single-objective MDP $(\mathcal{S},\mathcal{A},{\bm{P}},{\bm{r}}^{\top}{\bm{w}},\gamma)$, where we use the preference vector ${\bm{w}}$ to transform $M$-objective reward tensor ${\bm{r}}\in\mathbb{R}^{S\times A\times D}$ to single-objective reward $\in\mathbb{R}^{S\times A}$. By Lemma 12, $\pi_{1}$ and $\pi_{2}$ are optimal deterministic policies on $(\mathcal{S},\mathcal{A},{\bm{P}},{\bm{r}}^{\top}{\bm{w}},\gamma)$. By Lemma 13, the optimal policies on $(\mathcal{S},\mathcal{A},{\bm{P}},{\bm{r}}^{\top}{\bm{w}},\gamma)$ lies in the convex hull constructed by $\pi_{1}$ and $\pi_{2}$, i.e.,
@@
Given a finite MDP $(\mathcal{S},\mathcal{A},{\mathbf{P}},{\bm{r}},\gamma)$, where ${\mathbf{P}}$ is the transition probability kernel and ${\bm{r}}\geq 0$ is the reward function. Let $\pi_{1}$ and $\pi_{2}$ be two deterministic policies.
@@
Let ${\mathcal{M}}$ be the set of states that $\pi_{1}$ and $\pi_{2}$ differ on. By definition of distance between two deterministic policies, $M\coloneqq|{\mathbb{M}}|=d(\pi_{1},\pi_{2})$.
@@
where $s_{t}$ denotes the state of the process at time $t$. For any $\bar{s}\in{\mathcal{M}}$, we define the random variable ${\mathbf{H}}^{\pi}(\bar{s},{\mathcal{M}})=0$ as it takes 00 steps from $\bar{s}^{\prime}$ to ${\mathcal{M}}.$ Moreover, we define the total return of $d$-th objective before hitting ${\mathcal{M}}$ from state $s$ under policy $\pi$ as the sum of discounted rewards until the process reaches any state in ${\mathcal{M}}$. It is given by:
@@
Denote the expectation of the total return and discount before hitting $\bar{s}$ conditioned on avoiding all other states in ${\mathcal{M}}$ before reaching $\bar{s}$ as $\tilde{{\mathbf{V}}}^{\pi}_{d}\in\mathbb{R}^{S\times 1}$, $\tilde{{\mathbf{V}}}^{\pi}\in\mathbb{R}^{S\times D}$, and $\tilde{\Gamma}\in\mathbb{R}^{S\times M}$, respectively. For $s\in\mathcal{S}$, $s^{\prime}\in{\mathcal{M}}$, $d\in\{1,\cdots,D\}$,
$\tilde{{\mathbf{V}}}^{\pi}_{d}(s,{\mathcal{M}})\coloneqq$ $\mathbb{E}\left[\tilde{{\mathbf{R}}}^{\pi}_{d}(s,{\mathcal{M}})\right],$ (8)
@@
By the ergodicity assumption of MDP, starting from any state, the probability of arriving ${\mathcal{M}}$ is $1$. Recall ${\mathbf{V}}^{\pi}\in\mathbb{R}^{S\times D}$ is the value function with initial state distribution $\mu$, where $\mu$ is ignored for the notation simplicity. Moreover, we use ${\mathbf{V}}^{\pi}_{d}(s)$ and ${\mathbf{V}}^{\pi}_{d}({\mathcal{M}})$ to denote the $d$-th objective value function at state $s$ and at state set ${\mathcal{M}}$, respectively. By the definition of value function, ${\mathbf{V}}^{\pi}_{d}(s)$ can be written as
@@
Note that $\Gamma^{\pi}$ and $\tilde{{\mathbf{V}}}^{\pi}$ are the same for any policy in $\Pi_{N}(s_{1},{\mathcal{M}})$. Since we only consider policies within $\Pi_{N}(s_{1},{\mathcal{M}})$, for simplicity of notation, we omit the superscripts $\pi$ in $\Gamma^{\pi}$ and $\tilde{{\mathbf{V}}}^{\pi}$ in the following sections, as they remain constant for any policies within $\Pi_{N}(s_{1},{\mathcal{M}})$.
@@
(1) When $\mathrm{rank}(\Lambda)\leq M$, $\Lambda$ has full row rank. This case also implies $M\leq D$. By Theorem 3, the only solution to ${\mathbf{X}}\Lambda={\bm{0}}$ is ${\mathbf{X}}={\bm{0}}$. Therefore, Eq. 17 is equal to
@@
When $M=S$, $\left[\Gamma,{\bm{1}}_{S}\right]\in\mathbb{R}^{S\times(S+1)}$ has rank $S$, and When $M<S$, Since $\left[\Gamma,{\bm{1}}_{S}\right]$ has $S$ rows and rank at least $M$. Hence, $\left[\Gamma,{\bm{1}}_{S}\right]$ has $S$ rows and rank at least $M$ in both cases. Fix $k$. By the rank-nullity theorem Theorem 3 and Lemma 4, for $\bar{s}\in{\mathcal{M}}$, if the solution of ${\mathbf{P}}^{\pi_{2}}(\bar{s})$ exists, the dimension of the solution ${\mathbf{P}}^{\pi_{2}}(\bar{s})$ is at most $S-M$. On the other hand, if no solution to ${\mathbf{P}}^{\pi_{2}}(\bar{s})$ exists, the dimension of the solution space is 00. Therefore, the dimension of solution space of ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ is at most $M\times(S-M)$.
@@
Since valid probability transition matrix ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ requires that the row sums equal one, the ambient space ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ lies in has dimension $M\times(S-1)$. Since the solution space of ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ that satisfies Eq. 17 has dimension at most $M\times(S-M)+1$ with $\mathrm{rank}(\Lambda)>0$, while the ambient space where ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ resides has dimension $M\times(S-1)$, the dimension of the solution space is significantly lower than that of the full space. Therefore, the set of points where ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ satisfies Eq. 17 forms a subset of Lebesgue measure 0 in the $M\times(S-1)$-dimensional space.
@@
Combining with the dimension of ${\mathbb{G}}$, the dimension of solution space of ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ that satisfies Eq. 17 is at most $M\times(S-M)+M\times(M-\mathrm{rank}(\Lambda))+1=M\times(S-\mathrm{rank}(\Lambda))+1$. Since the solution space of ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ that satisfies Eq. 17 has dimension at most $M\times(S-\mathrm{rank}(\Lambda))+1$ with $\mathrm{rank}(\Lambda)>0$, while the ambient space where ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ resides has dimension $M\times(S-1)$, the dimension of the solution space is significantly lower than that of the full space. Therefore, according to results from measure theory, the set of points where ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ satisfies Eq. 17 forms a subset of Lebesgue measure 0 in the $M\times(S-1)$-dimensional space.

### v2

@@ 4.3 Locality Property of the Pareto front
Let deterministic policy $\pi$ correspond to a vertex on ${\mathbb{J}}$. We denote the set of neighbors on the boundary of ${\mathbb{J}}$ to $J^{\pi}$ as $\mathrm{N}(J^{\pi},{\mathbb{J}})$, and the set of deterministic policies whose long-term returns make up $\mathrm{N}(J^{\pi},{\mathbb{J}})$ as $\mathrm{N}_{\pi}(J^{\pi},{\mathbb{J}})$. Let ${\mathbb{F}}(J^{\pi},{\mathbb{J}})$ and ${\mathbb{F}}(J^{\pi},\mathrm{Conv}(J^{\Pi_{1}}\cup\{J^{\pi}\})$ be the set of incident faces of $J^{\pi}$ on ${\mathbb{J}}$ and $\mathrm{Conv}(J^{\Pi_{1}}\cup\{J^{\pi}\})$, respectively.
@@
The proof is provided in Theorem 6 and Proposition 3 of Appendix F. This theorem guarantees that we can find all neighboring vertices and faces on the boundary of ${\mathbb{J}}$ from the local convex hull $\mathrm{Conv}(J^{\Pi_{1}}\cup\{J^{\pi}\})$. By iterating through the process of constructing the convex hull of all distance-one policies and retaining the neighboring faces, we can eventually traverse the entire boundary of ${\mathbb{J}}$.
We also establish key properties of the Pareto front on a convex polytope (proof is provided in Appendix G).
@@
Given a convex polytope ${\mathbb{C}}$, let $F$ be face of ${\mathbb{C}}$ defined as the intersection of $n$ facets, i.e., $F=\cap_{i=1}^{n}F_{i}$, where $n\geq 1$ and each facet is of the form $F_{i}={\mathbb{C}}\cap\{{\bm{x}}\in\mathbb{R}^{d}:{\bm{w}}_{i}^{\top}{\bm{x}}=c_{i}\}$. Then the face $F$ lies on the Pareto front of ${\mathbb{C}}$ if and only if there exists a linear combination of the normal vectors, weighted by $\boldsymbol{\alpha}_{i}\geq 0$ for any $i\in\{1,\cdots,n\}$, such that $\sum_{i}\boldsymbol{\alpha}_{i}[{\bm{w}}_{i}]_{j}>0$ where $[{\bm{w}}_{i}]_{j}$ is the $j$-th component of ${\bm{w}}_{i}$.
Based on Lemma 3, we can also extract the Pareto optimal faces from the local convex hull. Let ${\mathbb{F}}(J^{\pi},\mathbb{P}({\mathbb{J}}))$ be the set of incident faces of $J^{\pi}$ on the Pareto front $\mathbb{P}({\mathbb{J}})$. The proof is shown in Proposition 6 of Appendix F.
@@ Appendix C Supporting Lemmas and Theorems
Let ${\bm{x}}$ be a point in a set ${\mathbb{S}}$. ${\bm{x}}$ lies on the Pareto front of ${\mathbb{S}}$ if and only if there exists a vector ${\bm{w}}>0$ such that ${\bm{w}}^{\top}({\bm{x}}-{\bm{y}})\geq 0$ for all ${\bm{y}}\in{\mathbb{S}}$.
@@
[…]
($\implies$) If ${\bm{x}}$ is Pareto optimal, then for any ${\bm{y}}\in{\mathbb{S}}$, ${\bm{x}}$ is at least as large as the corresponding objective of ${\bm{y}}$ and ${\bm{x}}$ is strictly better than ${\bm{y}}$ in at least one objective. Given ${\bm{w}}>0$, ${\bm{w}}^{\top}({\bm{x}}-{\bm{y}})\geq 0$. ∎
@@
[38] The occupancy measure of any policy $\pi$ and initial distribution $\mu$ is defined as $d_{\mu}^{\pi}(s)\coloneqq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s|s_{0}\sim\mu(\cdot))$. Similarly, define $\bar{d}_{\mu}^{\pi}(s,a)\coloneqq(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}P(s_{t}=s,a_{t}=a|s_{0}\sim\mu(\cdot))=d_{\mu}^{\pi}(s)\pi(a|s)$. Define for any set of policies $U$, $L^{U}(\mu)\coloneqq\cup_{\pi\in U}\{\bar{d}_{\mu}^{\pi}\}$ and define $L(\mu)\coloneqq\cup_{\pi\in\Pi}\{\bar{d}_{\mu}^{\pi}\}$. Then, $L(\mu)$ is closed convex polytope and $L(\mu)=\mathrm{Conv}(L^{\Pi_{D}}(\mu))$.
@@
Let $\Pi_{\text{all}}$ be the set of all policies, including both stationary and non-stationary policies, and let ${\mathbb{J}}_{\text{all}}(\mu)=\{J^{\pi}(\mu)|\pi\in\Pi_{\text{all}}\}$ denote the set of achievable long-term returns for all policies in $\Pi_{\text{all}}$. Recall that ${\mathbb{J}}(\mu)=\{J^{\pi}(\mu)|\pi\in\Pi\}$ represents the set of achievable long-term returns for all policies in the set of stationary policies $\Pi$. By Theorem 3.1 of [38], it is sufficient to represent the ${\mathbb{J}}(\mu)$ using only stationary policies, i.e., ${\mathbb{J}}_{\text{all}}(\mu)={\mathbb{J}}(\mu)$. Therefore, we focus on the set of stationary policies to construct the Pareto front in our paper.
The Pareto front of ${\mathbb{J}}(\mu)$ is denoted as $\mathbb{P}({\mathbb{J}}(\mu))$. Lemma 1 shows that the ${\mathbb{J}}(\mu)$ construct a convex polytope. The boundary and vertices of the convex polytope ${\mathbb{J}}(\mu)$ are denoted as ${\mathbb{B}}({\mathbb{J}}(\mu))$ and ${\mathbb{V}}({\mathbb{J}}(\mu))$, respectively. Lemma 1 also implies the vertices of ${\mathbb{J}}(\mu)$ are deterministic policies, that is, ${\mathbb{V}}({\mathbb{J}}(\mu))\subseteq\Pi_{D}$.
@@
(Restatement of Lemma 1) ${\mathbb{J}}(\mu)$ is a closed convex polytope, and the vertices of ${\mathbb{J}}(\mu)$ can be achieved by deterministic policies.
@@
Under 1, for any edge on ${\mathbb{B}}({\mathbb{J}}(\mu))$ with endpoints denoted as $J^{\pi_{1}}(\mu)$ and $J^{\pi_{2}}(\mu)$, any point on this edge can only be achieved by the long-term return of a convex combination of policies $\pi_{1}$ and $\pi_{2}$. Specifically, for any $\alpha\in[0,1]$, there exists a $\beta\in[0,1]$ such that
@@
We construct a single-objective MDP $(\mathcal{S},\mathcal{A},{\bm{P}},{\bm{r}}^{\top}{\bm{w}},\gamma)$, where we use the preference vector ${\bm{w}}$ to transform $M$-objective reward tensor ${\bm{r}}\in\mathbb{R}^{S\times A\times D}$ to single-objective reward $\in\mathbb{R}^{S\times A}$. Under 1, $\pi_{1}$ and $\pi_{2}$ are optimal deterministic policies on $(\mathcal{S},\mathcal{A},{\bm{P}},{\bm{r}}^{\top}{\bm{w}},\gamma)$ by Lemma 12. By Lemma 13, the optimal policies on $(\mathcal{S},\mathcal{A},{\bm{P}},{\bm{r}}^{\top}{\bm{w}},\gamma)$ lies in the convex hull constructed by $\pi_{1}$ and $\pi_{2}$, i.e.,
@@
Given a discounted finite MDP $(\mathcal{S},\mathcal{A},{\mathbf{P}},{\bm{r}},\gamma)$, where ${\mathbf{P}}$ is the transition probability kernel and ${\bm{r}}\geq 0$ is the reward function. Let $\pi_{1}$ and $\pi_{2}$ be two deterministic policies. Under 1, we have
@@
Let ${\mathcal{M}}$ be the set of states that $\pi_{1}$ and $\pi_{2}$ differ on. By definition of distance between two deterministic policies, $M\coloneqq|{\mathcal{M}}|=d(\pi_{1},\pi_{2})$.
Recall that $\Pi_{N}(\pi,{\mathcal{M}})$ denotes the set of deterministic policies that differ from $\pi$ only at states within ${\mathcal{M}}$, while matching $\pi$ at all other states. In other words, all policies in $\Pi_{N}(\pi,{\mathcal{M}})$ take the same actions for states outside the set ${\mathcal{M}}$. For states that cannot reach ${\mathcal{M}}$ within a finite number of steps, the value function is identical across all such policies. Conversely, for states that can reach ${\mathcal{M}}$ in finite steps, the value function depends on the policy within ${\mathcal{M}}$. With some slight abuse of notation, in the remainder of this proof, we focus only on the set of states that can reach ${\mathcal{M}}$ within a finite number of steps, denoted as ${\mathcal{S}}$.
@@
where $s_{t}$ denotes the state of the process at time $t$. For any $\bar{s}\in{\mathcal{M}}$, we define the random variable ${\mathbf{H}}^{\pi}(\bar{s},{\mathcal{M}})=0$ as it takes 00 steps from $\bar{s}$ to ${\mathcal{M}}.$ Moreover, we define the total return of $d$-th objective before hitting ${\mathcal{M}}$ from state $s$ under policy $\pi$ as the sum of discounted rewards until the process reaches any state in ${\mathcal{M}}$. It is given by:
@@
Denote the expectation of the total return and discount before hitting $\bar{s}$ conditioned on avoiding all other states in ${\mathcal{M}}$ before reaching $\bar{s}$ as $\tilde{{\mathbf{V}}}^{\pi}_{d}=[\tilde{V}^{\pi}_{d}(s_{0},{\mathcal{M}}),\cdots,\tilde{V}^{\pi}_{d}(s_{|\mathcal{S}|},{\mathcal{M}})]^{T}\in\mathbb{R}^{S\times 1}$, $\tilde{{\mathbf{V}}}^{\pi}\in\mathbb{R}^{S\times D}$, and $\tilde{\Gamma}\in\mathbb{R}^{S\times M}$, respectively. For $s\in\mathcal{S}$, $s^{\prime}\in{\mathcal{M}}$, $d\in\{1,\cdots,D\}$, we define
$\tilde{V}^{\pi}_{d}(s,{\mathcal{M}})\coloneqq$ $\mathbb{E}\left[\tilde{R}^{\pi}_{d}(s,{\mathcal{M}})\right],$ (8)
@@
Recall ${\mathbf{V}}^{\pi}\in\mathbb{R}^{S\times D}$ is the value function with initial state distribution $\mu$, where $\mu$ is ignored for the notation simplicity. Moreover, we use ${\mathbf{V}}^{\pi}_{d}(s)$ and ${\mathbf{V}}^{\pi}_{d}({\mathcal{M}})$ to denote the $d$-th objective value function at state $s$ and at state set ${\mathcal{M}}$, respectively. By the definition of value function, ${\mathbf{V}}^{\pi}_{d}(s)$ can be written as
@@
Note that $\Gamma^{\pi}$ and $\tilde{{\mathbf{V}}}^{\pi}$ are the same for any policy in $\Pi_{N}(\pi_{1},{\mathcal{M}})$. Since we only consider policies within $\Pi_{N}(\pi_{1},{\mathcal{M}})$, for simplicity of notation, we omit the superscripts $\pi$ in $\Gamma^{\pi}$ and $\tilde{{\mathbf{V}}}^{\pi}$ in the following sections, as they remain constant for any policies within $\Pi_{N}(\pi_{1},{\mathcal{M}})$.
@@
(1) When $\mathrm{rank}(\Lambda)=M$, $\Lambda$ has full row rank. This case also implies $M\leq D$. By Theorem 3, the only solution to ${\mathbf{X}}\Lambda={\bm{0}}$ is ${\mathbf{X}}={\bm{0}}$. Therefore, Eq. 17 is equal to
@@
When $M=S$, $\left[\Gamma,{\bm{1}}_{S}\right]\in\mathbb{R}^{S\times(S+1)}$ has rank $S$. When $M<S$, Since $\left[\Gamma,{\bm{1}}_{S}\right]$ has $S$ rows and rank at least $M$. Hence, $\left[\Gamma,{\bm{1}}_{S}\right]$ has $S$ rows and rank at least $M$ in both cases. Fix $k$. By the rank-nullity theorem Theorem 3 and Lemma 4, for $\bar{s}\in{\mathcal{M}}$, if the solution of ${\mathbf{P}}^{\pi_{2}}(\bar{s})$ exists, the dimension of the solution ${\mathbf{P}}^{\pi_{2}}(\bar{s})$ is at most $S-M$. On the other hand, if no solution to ${\mathbf{P}}^{\pi_{2}}(\bar{s})$ exists, the dimension of the solution space is 00. Therefore, the dimension of solution space of ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ is at most $M\times(S-M)$.
@@
Since valid probability transition matrix ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ requires that the row sums equal one, the ambient space ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ lies in has dimension $M\times(S-1)$. Since the solution space of ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ that satisfies Eq. 17 has dimension at most $M\times(S-M)+1$ with $\mathrm{rank}(\Lambda)>0$, while the ambient space where ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ resides has dimension $M\times(S-1)$. When $M>1$, the dimension of the solution space is significantly lower than that of the full space. Therefore, the set of points where ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ satisfies Eq. 17 forms a subset of Lebesgue measure 0 in the $M\times(S-1)$-dimensional space.
@@
Combining with the dimension of ${\mathbb{G}}$, the dimension of solution space of ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ that satisfies Eq. 17 is at most $M\times(S-M)+M\times(M-\mathrm{rank}(\Lambda))+1=M\times(S-\mathrm{rank}(\Lambda))+1$. Since the solution space of ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ that satisfies Eq. 17 has dimension at most $M\times(S-\mathrm{rank}(\Lambda))+1$ with $\mathrm{rank}(\Lambda)>0$, while the ambient space where ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ resides has dimension $M\times(S-1)$. When $M>1$, the dimension of the solution space is significantly lower than that of the full space. Therefore, according to results from measure theory, the set of points where ${\mathbf{P}}^{\pi_{2}}({\mathcal{M}})$ satisfies Eq. 17 forms a subset of Lebesgue measure 0 in the $M\times(S-1)$-dimensional space.